OpenStudy (anonymous):

What is the first step in solving this system of equations using elimination? 2x – 4y = 6

OpenStudy (solomonzelman):

there is only one equation here.

OpenStudy (xguardians):

The first step is to make opposites of one variable. IE: 2x + 5y = 10 x + 10 y = 50 Multiply the second equation by -2 to make them line up. 2x + 5y = 10 -2x - 20 y = -100 Then add the two together: -15 y = 90 But this doesn't relate to your problem since you only have one equation, you need two. a few moments ag